output matehing In this assignment you are to design and implement in ADS two networks: an LC matching network and a microstrPaper Design only!!
Answer #1

L L1 L-9.736031 nH R=1e-12 Ohm ormm Term Term Term1 Num 1 Z=25+j*50 Ohm Term2 Num-2 Z=50 Ohm C 1 C=4.113942 pF S-PARAMETERS SS-Parameters vs. Frequency Input Reflection Coefficient ADS ADS 10- 20 02 08 04 0.8 frea. GHz freq (0.0000Hz to 2.000GHZ) ADSMLSC TL4 Subst=MSub1 W 0.490204 mm L 10.596600 mn MSub MLIN TL3 MSUB Subst=MSub1 MSub1 W 0.490204 mm Term H 0.5 mm Term LS-Parameters vs. Frequency Input Reflection Coefficient ADS 30- 40- 20 02 14 0.8 freq, GHz freg (0.0000Hz to 2.000GHZ) ADS L0Using Madeling Lc L-9.736 nH Team2 Zond Ct 4.11PP Son 25t503 T71 アアプケ and impleented ploted. A The duign abans Circnit and th